# Stochastic Gravitational Wave Background from Highly-Eccentric Stellar-Mass Binaries in the Milli-hertz Band

### v1

@@ I Introduction
Various dynamic environments can give rise to stellar-mass bursting GW sources. For example, in dense star clusters, compact object binaries may attain non-negligible eccentricity due to external perturbations like GW capture, binary-single, and binary-binary scattering [e.g., 35, 36, 37, 38, 39, 4, 40, 41, 42, 43, 44, 45, 46, 47]. Highly eccentric mergers can also arise in stellar disks or active galactic nucleus accretion disks [48, 16, 49, 50]. Additionally, fly-by interactions and galactic tides may excite the eccentricity of wide compact object binaries in the galactic field [e.g., 51, 52, 53], potentially leading to observable mergers. Moreover, in a hierarchical triple system (a tight binary orbiting a third body on a much wider ”outer orbit”), the inner binary can undergo eccentricity oscillations via the eccentric Kozai-Lidov (EKL) mechanism [54, 55, 56], becoming a GW source with high eccentricity in the LISA band. These channels can significantly contribute to the overall merger rate of stellar-mass compact objects [e.g., 57, 58, 59, 60, 41, 61, 62, 63, 9].
@@
In this paper, our focus is on the stochastic GW background (GWB) originating from these bursting sources. In particular, as a collective GW signal of unresolved sources, the astrophysical stochastic GWB will compete with other expected LISA science sources, potentially producing a confusion noise that affects LISA sensitivity. Previous efforts have aimed to characterize various stochastic backgrounds, such as the collective GW signal from the galactic population of double white dwarf (DWD) binaries [see, e.g., 64, 65, 66, 67], extragalactic sources [20, Bonetti+2020Enoki2007, 68, 69, 70], and cosmological stochastic backgrounds [71, 72, 73, 74, 40, 75]. Additionally, it was recently suggested that extreme mass ratio inspirals (EMRIs) might have orders of magnitude higher formation rate [76] that may yield an order of magnitude noise level above LISA’s sensitivity level [77]. By understanding these backgrounds, we not only address the potential impact of confusion noise on the parameter estimation of other resolved sources but also extract information about the stochastic background itself, thus constraining the population of GW sources [see, e.g., 78, 77].
@@
As a natural consequence of dynamical formation, the highly eccentric, stellar-mass binaries may create a significant GW background in the LISA band. For example, in our previous studies [33], we anticipate the presence of approximately 3 to 45 detectable bursting binary black holes (BBHs) within the Milky Way, each with a signal-to-noise ratio (SNR) exceeding 5 for the upcoming LISA mission [33]. Beyond these detectable cases, there could be a considerably larger number of bursting sources with SNR values falling below the detection threshold, yet their collective contribution remains significant.
@@
Furthermore, the transient nature of these sources poses challenges in extracting their signals from the detector’s output [79, 31, 32], which potentially leads to a higher noise background level. In particular, the burst detection methods for stellar-mass binaries, especially when multiple sources’ bursts are present in the data stream simultaneously, remain relatively underdeveloped. Thus, there is uncertainty regarding our ability to identify all bursting sources in the detector’s output, even if the sources’ SNR is above the detection threshold. Moreover, the astrophysical burst signals may be intertwined with instrumental noise, such as glitches [see, e.g., 80, 81, 82], and contribute to the unresolved GW background.
@@
We note that, there have been previous works examining the GWB from circular stellar-mass BBHs [see, e.g., 73, 74, 83] and eccentric stellar-mass BBHs at a cosmological distance [see, e.g., 84, 40, 85]. These studies suggested that stellar-mass BBHs can make a non-negligible contribution to future observations, with eccentricity potentially affecting the overall shape of the millihertz GW background. In this work, however, we extend this analysis beyond the cosmological population of BBHs to explore the GWB contribution of highly eccentric sources at close distances, specifically bursting sources in the Milky Way and nearby galaxies. As discussed before, even considering only the Milky Way population, the number of these sources may be substantial [33], and their close distance can potentially yield a significant overall GWB level. Furthermore, the transient nature of their GW emission can affect the pattern of the GWB, with a single bursting source capable of emitting GWs in a wide range of frequencies, which may cause confusion in future observations. In our analysis, we consider the realistic spatial distribution and non-equilibrium formation history of highly eccentric sources.
@@
We organize this paper as follows. In Section II.1, we first demonstrate the properties of bursting GW sources, then discuss their detectability (Section II.2.1) and GW background calculation (Section II.2.2). In Section III, we introduce the population model of bursting sources, then carry out numerical simulations to predict the GW background level from highly-eccentric, stellar-mass BBHs in the Universe (Section III.2). We show the results separately, for the globular clusters (Section III.2.1), galactic field (Section III.2.2), and the galactic nucleus (Section III.2.3). In Section IV, we summarize and discuss the overall effect of stochastic GW background from stellar-mass bursting sources.
@@ II.1 The Properties of GW Bursts from Highly Eccentric Binaries
$h_{\rm burst}$ $\sim\sqrt{\frac{32}{5}}\frac{m_{1}m_{2}}{D_{l}a(1-e)}$
$\sim 7.65\times 10^{-21}\eta_{s}\left(\frac{m_{\rm bin}}{20\rm M_{\odot}}\right)^{\frac{5}{3}}\left(\frac{f_{\rm burst}}{3.16\rm mHz}\right)^{\frac{2}{3}}\left(\frac{D_{l}}{8\rm kpc}\right)^{-1}\ ,$ (3)
in which $f_{\rm orb}=1/T_{\rm orb}$ is the orbital frequency of the bursting source, $D_{l}$ is the luminosity distance of the binary, and $\eta_{s}=4m_{1}m_{2}/(m_{1}+m_{2})^{2}$ is unity for equal mass sources.
@@
here $T_{\rm obs}$ is the observational time of the GW detector, and $S_{\mathrm{n}}(f)$ is the spectral noise density of LISA evaluated at GW frequency $f$ [see e.g., 88, 89]
@@
$\left(\frac{D_{l}}{8\rm kpc}\right)^{-1}\left(\frac{T_{\rm obs}}{4\rm yr}\right)^{\frac{1}{2}}\left(\frac{S_{n}(f_{\rm burst})}{S_{n}(3.16\rm mHz)}\right)^{-\frac{1}{2}}.$ (7)
@@
where $\mu=m_{1}m_{2}/(m_{1}+m_{2})$, and $q=m_{1}/m_{2}$.
@@
We note that Equation (10) is based on the comparison in the same GW frequency band and does not take into account the sources’ detectability. However, since circular inspirals have a larger SNR than the highly eccentric configuration, they may exceed the LISA detection threshold below the millihertz band, extending their detectable time and enlarging the detected population. For example, in our previous work [33], we estimated the detectability of a $10-10{\rm M_{\odot}}$ BBH system with $e\sim 0,a\sim 0.03$ au in the Milky Way ($\sim 10\rm kpc$). It turns out that such a system will be detectable for $10^{7}$ yr, with an orbital period of $T_{\rm orb}\sim 10^{4}$ s and GW frequency $f_{\rm circ,GW}\sim 10^{-4}$ Hz. Because we expect a large population of circular DWDs in the sub-millihertz band [see, e.g., 90, 67], the identification of the near-circular, low-frequency systems is beyond the scope of this study. Thus, in this paper, we focus on the dynamically formed, highly-eccentric bursting sources and limit our aforementioned argument of the population in the same frequency band (e.g., $f_{\rm GW}\gtrsim 1\rm mHz$).
@@ II.2.1 Identification of Bursting Signals and Residual Noise
Bursting sources have the potential to form a non-negligible GW background in the millihertz band, given their significant GW radiation (Equations (2)-(7)) and extended lifetime (Equation (8)). Particularly, frequent bursts from multiple sources with low SNR can accumulate to yield a stochastic background noise (see Eqs. (6) and (7)). For example, in Figure 1, we follow the strategy described in Section III and calculate the collective GW signal from bursting BBHs in the Milky Way globular clusters. As shown in this example, the collective bursting signal is a superposition of GW transients, i.e., individual bursts, from various sources, each localized in time and frequency. Additionally, we also include the LISA detector’s response to the bursting GW background, which yields a long-term modulation due to the orbit of LISA around the Sun (see Upper Panel of Figure 1).
The transient nature of bursting signals (see Lower Panel of Figure 1) poses several challenges for existing data analysis strategies, such as the matched filtering method [91, 34], which relies on accurate templates of GW signals and may not perform optimally in fitting bursts [see, e.g., 79, 31, 32]. In particular, GW templates for highly eccentric sources are underdeveloped, which adds to the difficulty of constructing a template bank for the extraction of the signal. Moreover, multiple bursting sources can be present in the detector’s output, which may lead to the misidentification of sources and degeneracy of the fitted parameters.
We note that many efforts have been made to analyze transient events in LIGO data analysis (e.g., power stacking [1], wavelet decomposition [92], and the Q-transform [93, 94]). Furthermore, several studies focused on the transients from highly eccentric EMRIs, which can be possibly seen by LISA [20, 95, 96, 21, 97]. However, the burst detection methods for stellar-mass binaries, particularly when multiple sources’ bursts are present in the datastream simultaneously, remain relatively underdeveloped. In other words, there is uncertainty regarding our ability to identify all bursting sources, even when their signal-to-noise ratio is sufficiently high.
Figure 1: GW background from bursting BBHs within the Milky Way globular clusters. Here we generate the bursting BBHs population using the methodology described in Section III, and calculate their collective GW singals observed by LISA. Upper panel shows the detector’s Michelson response [99, 100], $s_{1}(t)$, over a 1-yr period. Lower panel shows a zoom-in of the signal. In the upper panel, the modulation of the GW background is a result of the detector’s annual motion around the sun.
@@
II.2.2 Stochastic Background Calculation
@@
For an eccentric binary, its time domain waveform, $h(a,e,t)$, can be decomposed into different harmonics, each with the frequency of $f_{n}=nf_{\rm{orb}}$ [for simplicity, here we rms averaged over the binary inclination, see e.g., 101, 102, 38]:
$h(a,e,t)=\sum_{n=1}^{\infty}h_{n}\left(a,e,f_{n}\right)\exp\left(2\pi if_{n}t\right),$ (12)
@@
$h_{n}\left(a,e,f_{n}\right)=\frac{2}{n}\sqrt{g(n,e)}h_{0}(a)\ ,$ (13)
and:
$h_{0}(a)=\sqrt{\frac{32}{5}}\frac{m_{1}m_{2}}{D_{l}a}\ ,$ (14)
@@
$\left.+\left(1-e^{2}\right)\left(J_{n-1}-2J_{n}+J_{n+2}\right)^{2}+\frac{4}{3n^{2}}J_{n}^{2}\right],$ (15)
in which $J_{i}$ is the i-th Bessel function evaluated at $ne$.
The energy density of a monochromatic GW signal with frequency $f$ can be expressed as [see, e.g., 103]:
@@
where $h_{+},h_{\times}$ are the amplitude of GW’s two polarizations, and $h_{0}$ stands for the sky-averaged stain amplitude.
Since the gravitational wave signal from an eccentric source consists of multiple harmonics with different frequencies, the total energy spectral density is a superposition of each harmonic’s contribution. In other words, by combining Equations (12)-(16), we can derive the energy density contribution from all the harmonics below a given frequency, $f_{c}=n_{\rm max}f_{\rm orb}$:
$\rho_{\rm ecc}(f_{c})=\sum_{n=1}^{n_{\rm max}}\frac{\pi}{4}\left(nf_{\rm orb}\right)^{2}\left[\frac{2}{n}\sqrt{g(n,e)}h_{0}(a)\right]^{2}\ .$ (17)
Here, we assume the GW source evolves slowly. Thus, the signal does not undergo a significant frequency shift during the observation, and each harmonic is a monochromatic signal with the power concentrates on $f=nf_{\rm orb}$. We expect that most sources will be consistent with this assumption [e.g., 63, 33]; however, in some cases, the signal may shift in a short timescale, [24, 62], which is beyond the scope of this study.
A key term in the noise spectral density, Eq. (11), is $d\rho/df$. However, as can be seen from Equation (12)-(17), the number of harmonics making a substantial contribution to the signal rises rapidly as eccentricity reaches extreme values. For example, we may have to consider millions to billions of harmonics when calculating the spectral density of GW signal from a binary with $e\gtrsim 0.95$. Consequently, Equation (17) might not always be practical for noise spectral density calculations.
Thus, we further average Equation (17) over a frequency bin $df$ to get a smoothed expression of $\rho(f)$. In particular, for a frequency bin $df$ at frequency $f$, each GW harmonic within this bin contributes an energy density of approximately $\sim\frac{\pi}{4}f^{2}\left[\frac{2}{n}\sqrt{g(n,e)}h_{0}(a)\right]^{2}$, where $n\approx f/f_{\rm orb}$, and the number of harmonics included in this bin equals $\sim df/f_{\rm orb}$. Therefore, the total energy spectral density at $f$ is obtained by summing the energy densities of all these harmonics:
$d\rho_{\rm{ecc,avg}}(f)=\frac{\pi}{4}f^{2}\left[\frac{2}{n}\sqrt{g(n,e)}h_{0}(a)\right]^{2}\times\frac{df}{f_{\rm orb}},$ (18)
@@
$\frac{d\rho_{\rm ecc,avg}}{df}=\frac{\pi}{4}f^{2}\left[\frac{2}{n}\sqrt{g(n,e)}h_{0}(a)\right]^{2}\frac{1}{f_{\rm orb}},$ (19)
@@
Below, we adopt Equation (19) to calculate the energy spectrum of a single bursting source, then sum all the bursting sources’ contributions to get the total GW power spectrum:
$S_{n}(f)=\frac{4}{\pi f^{2}}\frac{d\rho}{df}=\sum_{\rm sources}\left[\frac{2}{n(f)}\sqrt{g(n(f),e)}h_{0}(a)\right]^{2}\frac{1}{f_{\rm orb}}\ .$ (20)
For bursting sources at a cosmological distance, we further take into account the redshift $z$ [e.g., see 77, 71]:

### v2

@@ I Introduction
Various dynamic environments can give rise to stellar-mass bursting GW sources. For example, in dense star clusters, compact object binaries may attain non-negligible eccentricity due to external perturbations like GW capture, binary-single, and binary-binary scattering [e.g., 35, 36, 37, 38, 39, 4, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Highly eccentric mergers can also arise in stellar disks or active galactic nucleus accretion disks [49, 16, 50, 51]. Additionally, fly-by interactions and galactic tides may excite the eccentricity of wide compact object binaries in the galactic field [e.g., 52, 53, 54], potentially leading to observable mergers. Moreover, in a hierarchical triple system (a tight binary orbiting a third body on a much wider ”outer orbit”), the inner binary can undergo eccentricity oscillations via the eccentric Kozai-Lidov (EKL) mechanism [55, 56, 57], becoming a GW source with high eccentricity in the LISA band. These channels can significantly contribute to the overall merger rate of stellar-mass compact objects [e.g., 58, 59, 60, 61, 41, 62, 63, 64, 9].
@@
In this paper, our focus is on the stochastic GW background (GWB) originating from these bursting sources. In particular, as a collective GW signal of unresolved sources, the astrophysical stochastic GWB will compete with other expected LISA science sources, potentially producing a confusion noise that affects LISA sensitivity. Previous efforts have aimed to characterize various stochastic backgrounds, such as the collective GW signal from the galactic population of double white dwarf (DWD) binaries [see, e.g., 65, 66, 67, 68], extragalactic sources [20, 69, 70, 71, 72, 73], and cosmological stochastic backgrounds [74, 75, 76, 77, 40, 78, 79]. Additionally, it was recently suggested that extreme mass ratio inspirals (EMRIs) might have orders of magnitude higher formation rate [80] that may yield an order of magnitude noise level above LISA’s sensitivity level [81]. By understanding these backgrounds, we not only address the potential impact of confusion noise on the parameter estimation of other resolved sources but also extract information about the stochastic background itself, thus constraining the population of GW sources [see, e.g., 82, 81].
@@
As a natural consequence of dynamical formation, the highly eccentric, stellar mass binaries may create a significant GW background in the LISA band. For example, in our previous studies [33], we anticipate the presence of approximately 3 to 45 detectable bursting binary black holes (BBHs) within the Milky Way, each with a signal-to-noise ratio (SNR) exceeding 5 for the upcoming LISA mission [33]. Beyond these detectable cases, there could be a considerably larger number of bursting sources with SNR values falling below the detection threshold, yet their collective contribution remains significant.
@@
Furthermore, the transient nature of these sources poses challenges in extracting their signals from the detector’s output [83, 31, 32], which potentially leads to a higher noise background level. In particular, the burst detection methods for stellar mass binaries, especially when multiple sources’ bursts are present in the data stream simultaneously, remain relatively underdeveloped. Thus, there is uncertainty regarding our ability to identify all bursting sources in the detector’s output, even if the sources’ SNR is above the detection threshold. Moreover, the astrophysical burst signals may be intertwined with instrumental noise, such as glitches [see, e.g., 84, 85, 86], and contribute to the unresolved GW background.
@@
We note that there have been previous works examining the GWB from circular stellar mass BBHs [see, e.g., 76, 77, 87] and eccentric stellar mass BBHs at a cosmological distance [see, e.g., 88, 40, 89]. These studies suggested that stellar-mass BBHs can make a non-negligible contribution to future observations, with eccentricity potentially affecting the overall shape of the millihertz GW background. In this work, however, we extend this analysis beyond the cosmological population of BBHs to explore the GWB contribution of highly eccentric sources at close distances, specifically bursting sources in the Milky Way and nearby galaxies. As discussed before, even considering only the Milky Way population, the number of these sources may be substantial [33], and their close distance can potentially yield a significant overall GWB level. Furthermore, the transient nature of their GW emission can affect the pattern of the GWB, with a single bursting source capable of emitting GWs in a wide range of frequencies, which may cause confusion in future observations. In our analysis, we consider the realistic spatial distribution and non-equilibrium formation history of highly eccentric sources.
@@
We organize this paper as follows. In Section II.1, we first demonstrate the properties of bursting GW sources, then discuss their detectability (Section II.2.1) and GW background calculation (Section II.3). In Section III, we introduce the population model of bursting sources, then carry out numerical simulations to predict the GW background level from highly-eccentric, stellar-mass BBHs in the Universe (Section III.2). We show the results separately, for the globular clusters (Section III.2.1), galactic field (Section III.2.2), and the galactic nucleus (Section III.2.3). In Section IV, we summarize and discuss the overall effect of stochastic GW background from stellar-mass bursting sources.
@@ II.1 The Properties of GW Bursts from Highly Eccentric Binaries
$h_{\rm burst}$ $\sim\sqrt{\frac{32}{5}}\frac{m_{1}m_{2}}{r_{d}a(1-e)}$
$\sim 7.65\times 10^{-21}\eta_{s}\left(\frac{m_{\rm bin}}{20\rm M_{\odot}}\right)^{\frac{5}{3}}\left(\frac{f_{\rm burst}}{3.16\rm mHz}\right)^{\frac{2}{3}}\left(\frac{r_{d}}{8\rm kpc}\right)^{-1}\ ,$ (3)
in which $f_{\rm orb}=1/T_{\rm orb}$ is the orbital frequency of the bursting source, $r_{d}$ is the comoving distance of the binary, and $\eta_{s}=4m_{1}m_{2}/(m_{1}+m_{2})^{2}$ is unity for equal mass sources.
@@
here $T_{\rm obs}$ is the observational time of the GW detector, and $S_{\mathrm{n}}(f)$ is the spectral noise density of LISA evaluated at GW frequency $f$ [see e.g., 92, 93]
@@
$\left(\frac{r_{d}}{8\rm kpc}\right)^{-1}\left(\frac{T_{\rm obs}}{4\rm yr}\right)^{\frac{1}{2}}\left(\frac{S_{n}(f_{\rm burst})}{S_{n}(3.16\rm mHz)}\right)^{-\frac{1}{2}}.$ (7)
@@
where $\mu=m_{1}m_{2}/(m_{1}+m_{2})$, and $q=m_{1}/m_{2}$. Note that this timescale is still much shorter than the Hubble time.
@@
We note that Equation (10) is based on the comparison in the same GW frequency band and does not take into account the sources’ detectability. However, since circular inspirals have a larger SNR than the highly eccentric configuration, they may exceed the LISA detection threshold below the millihertz band, extending their detectable time and enlarging the detected population. For example, in our previous work [33], we estimated the detectability of a $10-10{\rm M_{\odot}}$ BBH system with $e\sim 0,a\sim 0.03$ au in the Milky Way ($\sim 10\rm kpc$). It turns out that such a system will be detectable for $10^{7}$ yr, with an orbital period of $T_{\rm orb}\sim 10^{4}$ s and GW frequency $f_{\rm circ,GW}\sim 10^{-4}$ Hz. Because we expect a large population of circular DWDs in the sub-millihertz band [see, e.g., 94, 68], the identification of the near-circular, low-frequency systems is beyond the scope of this study. Thus, in this paper, we focus on the dynamically formed, highly-eccentric bursting sources and limit our aforementioned argument of the population in the same frequency band (e.g., $f_{\rm GW}\gtrsim 1\rm mHz$).
@@ II.2.1 Identification of Bursting Signals and Residual Noise
Bursting sources have the potential to form a non-negligible GW background in the millihertz band, given their significant GW radiation (Equations (2)-(7)) and extended lifetime (Equation (8)). Particularly, frequent bursts from multiple sources with low SNR can accumulate to yield a stochastic background noise (see Eqs. (6) and (7)). For example, in Figure 1, we show the simulation result of the collective GW signal from bursting BBHs in the Milky Way globular clusters. We randomly generate BBH systems and estimate their ages based on a constant formation rate $\gamma_{\rm GC}=10^{-6}\,\mathrm{yr}^{-1}$ in the Milky Way (for details of the population model, see Section III). At the time of observation, we numerically compute the GW signal of each individual system (excluding merged ones) using the x-model [95], then combine them to obtain the collective GW background in the time domain. Additionally, we also include the LISA detector’s response to the bursting GW background, which yields a long-term modulation due to the orbit of LISA around the Sun (see Upper Panel of Figure 1). As shown in this example, the collective bursting signal is a superposition of GW transients, i.e., individual bursts, from various sources, each localized in time and frequency.
The transient nature of bursting signals (see Lower Panel of Figure 1) poses several challenges for existing data analysis strategies, such as the matched filtering method [96, 34], which relies on accurate templates of GW signals and may not perform optimally in fitting bursts [see, e.g., 83, 31, 32]. In particular, GW templates for highly eccentric sources are underdeveloped, which adds to the difficulty of constructing a template bank for the extraction of the signal. Moreover, multiple bursting sources can be present in the detector’s output, which may lead to the misidentification of sources and degeneracy of the fitted parameters.
We note that many efforts have been made to analyze transient events in LIGO data analysis (e.g., power stacking [1], wavelet decomposition [97], and the Q-transform [98, 99]). Furthermore, several studies focused on the transients from highly eccentric EMRIs, which can be possibly seen by LISA [20, 100, 101, 21, 102]. However, the burst detection methods for stellar mass binaries, particularly when multiple sources’ bursts are present in the datastream simultaneously, remain relatively underdeveloped. In other words, there is uncertainty regarding our ability to identify all bursting sources, even when their signal-to-noise ratio is sufficiently high.
Figure 1: GW background from bursting BBHs within the Milky Way globular clusters. Here we generate the bursting BBHs population using the methodology described in Section III, and calculate their collective GW signals observed by LISA. Upper panel shows the detector’s Michelson response [104, 105], $s_{1}(t)$, over a 1-yr period. Lower panel shows a zoom-in of the signal. In the upper panel, the modulation of the GW background is a result of the detector’s annual motion around the sun.
@@
II.3 Stochastic Background Calculation
II.3.1 Stochastic Background from GW Sources with Fixed Orbital Parameters
@@
For a single eccentric binary, its time domain waveform, $h(a,e,t)$, can be decomposed into different harmonics, each with the frequency of $f_{n}$ [for simplicity, here we rms averaged over the binary inclination, see e.g., 106, 107, 38, 108]:
$h(a,e,t)=\sum_{n=1}^{\infty}h_{\mathrm{c},n}\left(a,e\right)\exp\left(2\pi if_{n}t\right),$ (12)
@@
$f_{n}=nf_{\rm orb}\,,\quad h_{\mathrm{c},n}=\frac{2}{n}\sqrt{g(n,e)}h_{0}(a)$ (13)
and
$h_{0}(a)=\sqrt{\frac{32}{5}}\frac{m_{1}m_{2}}{r_{d}a}\ ,$ (14)
@@
$\left.+\left(1-e^{2}\right)\left(J_{n-2}-2J_{n}+J_{n+2}\right)^{2}+\frac{4}{3n^{2}}J_{n}^{2}\right],$ (15)
in which $J_{i}$ is the i-th Bessel function evaluated at $ne$. Note that $r_{d}$ is the comoving radial distance to the binary.
The energy density of a monochromatic GW signal with frequency $f$ can be expressed as [see, e.g., 109]:
@@
where $h_{+},h_{\times}$ are the amplitude of GW’s two polarizations, and $h_{0}$ stands for the sky-averaged stain amplitude in Equation (14).
Since the gravitational wave signal from an eccentric source consists of multiple harmonics with different frequencies, the energy density is a superposition of each harmonic’s contribution. In other words, by combining Equations (12)-(16), we can sum the contributions of all harmonics to determine the total gravitational wave energy density of an eccentric binary system:
$\rho_{\rm ecc}=\sum_{n=1}^{\infty}\frac{\pi}{4}f_{n}^{2}h_{\mathrm{c},n}^{2}\ .$ (17)
Here, we assume the GW source evolves slowly. Thus, the signal does not undergo a significant frequency shift during the observation, and each harmonic is a monochromatic signal where the power concentrates on $f=nf_{\rm orb}$ (see Equation (16)). We expect that most sources will be consistent with this assumption for the harmonics that contribute most to the detection signal-to-noise ratio [e.g., 35, 64, 33]; however, in some cases, the signal may shift in a short timescale, [24, 63], which is beyond the scope of this study.
A key term in the noise spectral density, Eq. (11), is $d\rho/df$. To get this term from Equation (17), we need to calculate the distribution of energy density as the function of the GW signal’s frequency. In other words, we sum the energy density contributions from the harmonics within a given frequency bin $(f_{i},f_{f})$:
[…]
Here, $n=n_{i},n_{i}+1,...\,n_{f}$ represents all the harmonic numbers between the initial and final frequencies, $f_{i}/f_{\rm orb}$ and $f_{f}/f_{\rm orb}$. Furthermore, as can be seen from Equation (12)-(18), the number of harmonics making a substantial contribution to the signal rises rapidly as eccentricity reaches extreme values. For example, we may have to consider millions to billions of harmonics when calculating the spectral density of GW signal from a binary with $e\gtrsim 0.95$. Consequently, using the sum in Equation (18) might not always be practical in noise spectral density calculations. Thus, for the highly eccentric systems in this paper, we average Equation (18) over the frequency bin to get a smoothed expression of $\rho(f)$. In particular, consider a frequency bin with the width of $df$ at frequency $f$. Each GW harmonic in this bin contributes an energy density of approximately $\sim\frac{\pi}{4}f^{2}h_{\mathrm{c},n(f)}^{2}$, where $n=f/f_{\rm orb}$, and the number of harmonics included in this bin equals $dn=df/f_{\rm orb}=ndf/f$. Therefore, the energy density in this bin for a single source, $d\rho_{\rm{ecc,single}}$, is obtained by multiplying the energy density of a single harmonic by the number of harmonics:
$d\rho_{\rm{ecc,single}}(f)=\frac{\pi}{4}f^{2}h_{\mathrm{c},n(f)}^{2}\times\frac{n(f)df}{f},$ (19)
@@
$\frac{d\rho_{\rm ecc,single}}{df}=\frac{\pi}{4}fh_{\mathrm{c},n(f)}^{2}n(f).$ (20)
@@
Below, we substitute Equation (20) into (11) to calculate the energy spectrum of a single bursting source, then sum all the bursting sources’ contributions to get the total GW power spectrum:
$S_{n,\rm loc}(f)=\frac{4}{\pi f^{2}}\frac{d\rho}{df}=\sum_{i\in\rm sources}\frac{n_{i}(f)h_{\mathrm{c},n_{i}(f)}^{2}}{f}\ ,$ (21)
where $n_{i}(f)=f/f_{{\rm orb},i}$ for the $i^{\rm th}$ source. Equation (21) is useful when calculating the bursting GW background from a group of local sources with determined orbital parameters (e.g., bursting BBHs in the Milky Way galactic center, at a given time in the numerical simulation).
II.3.2 Time-Averaged Stochastic Background Under the Steady State Approximation
For some bursting BBH populations, we assume a continuous birth and death of compact object binaries while keeping the total number of GW sources unchanged (i.e., adopt the steady-state approximation). Therefore, to obtain the expectation value of total GW power spectral density, we further average Equation (21) over the evolution of highly eccentric systems.
In particular, first, consider the contribution of the $i^{\rm th}$ bursting source in Equation (21). The average of this contribution over the single system’s time evolution is:
$\langle S_{n,\rm loc}(f)\rangle_{\rm single,i^{\rm th}}=\int_{t_{0}}^{t_{0}+\tau_{\rm merger}}\frac{h_{\mathrm{c},n_{i}(f)}^{2}(f,t)}{f_{{\rm orb},i}(t)}\frac{dt}{\tau_{\rm merger}}\ ,$ (22)
in which $h_{\mathrm{c},n}(f,t)=\frac{2}{n(f,t)}\sqrt{g(n(f,t),e(t))}h_{0}(a(t))$ (see Equation (13)), and $n(f,t)=f/f_{\rm orb}(t)$.
In Equation (22), $t_{0}$ is the formation time of the BBH system in the simulations, and $\tau_{\rm merger}$ represents the merger time of the BBH system with initial orbital parameters $a(t_{0}),e(t_{0})$. When considering the long-term evolution of the source, parameters such as orbital frequency ($f_{\rm orb}$), semi-major axis ($a$), and eccentricity ($e$) become functions of time. For bursting systems undergoing isolated evolution, these parameters can be determined using Eqs. 5.6 - 5.14 in Peters [17]. However, for bursting systems affected by dynamical effects, such as the eccentric Kozai-Lidov mechanism, numerical simulations are required to determine the merger time and the evolution of other orbital parameters.
When integrating Equation (22) over the lifetime of a GW source, we determine its time-averaged GW power spectral density. Under the steady-state approximation, where the ensemble average equals the time average, Equation (22) provides the expectation value of spectrum contribution from the $i^{\rm th}$ GW source in observation.
Furthermore, we average Equation (22) over different BBH systems’ initial conditions in the population synthesis. This average is computed by summing the GW power spectral density of all the simulated systems, $\langle S_{n,\rm loc}(f)\rangle_{\rm single,i^{\rm th}}$ (see Equation (22)), and dividing by the number of systems generated in the simulation, $N_{s}$:
$\langle S_{n,\rm loc}(f)\rangle\bigg{|}_{\rm single,avg}=\frac{1}{N_{s}}\sum_{\rm i=1}^{N_{s}}\langle S_{n,\rm loc}(f)\rangle_{\rm single,i^{\rm th}}\ .$ (23)
We note that $N_{s}$ is artificially introduced to enlarge the sample size and ensure numerical accuracy. It differs from the actual number of systems observed.
Finally, we multiply Equation (23) with the expected number of bursting BBHs in observation, $\langle N_{\rm tot}\rangle$, and get the collective GW power spectrum:
$\langle S_{n,\rm loc}(f)\rangle\bigg{|}_{\rm tot}=\frac{\langle N_{\rm tot}\rangle}{N_{s}}\sum_{\rm i=1}^{N_{s}}\langle S_{n}(f)\rangle_{\rm single,i}\ .$ (24)
Here, $\langle N_{\rm tot}\rangle=\langle\tau_{\rm merger}\rangle\gamma_{\rm merger}$, where $\langle\tau_{\rm merger}\rangle$ represents the average lifetime of merger systems in the simulation, and $\gamma_{\rm merger}=dN/dt$ denotes the BBH merger rate per unit time in the considered region (see Section III.1). Equation (24) is useful when calculating the expectation of bursting GW background from local sources in the Monte Carlo simulation (e.g., bursting BBHs in the Milky Way globular clusters, assuming they have reached equilibrium distribution with a fixed merger rate.).
II.3.3 Average Stochastic Background from Extragalactic Sources
For bursting sources at a cosmological distance, we further take into account the redshift $z$ [e.g., see 81, 74]: